OpenStudy (anonymous):

Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

OpenStudy (anonymous):

y = −x2 + 4 y = 2x + 1

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

just need the steps

OpenStudy (anonymous):

there are many diffrent ways to fo this but we r going to do this this way, staep 1 take x and y and put them on 1 side

OpenStudy (anonymous):

step*

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

y +x^2= 4 y - 2x = 1 then subtract y from y

OpenStudy (anonymous):

i think i dont i just oearned his last month

OpenStudy (anonymous):

yeah i dont no sorry

OpenStudy (anonymous):

thats a hard question

OpenStudy (anonymous):

1. Subtract the 2 equations and factor them out and set equal to 0 to solve for x: \[-x^2-2x+3=0\rightarrow -(x^2+2x-3)=0\rightarrow -[(x+3)(x-1)]=0\] 2. Then, solve for x: \[x = -3 and 1\] 3. Plug x in to solve for y for 1 solution: \[y = 2x+1\rightarrow y = 2(1)+1\rightarrow y = 3\] 4. Plug in x to solve y for 2nd solution: \[y = 2x + 1\rightarrow y = 2(-3)+1\rightarrow y = -5\] 5. Solutions are (1,3) and (-3,-5)